OpenStudy (anonymous):

An electronics store makes a profit of $72 for every television sold and $90 for every computer sold. The manager's target is to make at least $360 a day on sales from televisions and computers. What are three possible solutions to the problem? A.(5,2), (3,3),and (1,4) are three possible solutions. B.(4,0), (2,2),and (1,1) are three possible solutions. C.(3,1), (2,2),and (1,0) are three possible solutions. D.(4,0), (3,3),and (1,4)are three possible solutions.

OpenStudy (anonymous):

3 TV's & 2 computers: 3*$72 + 2*$90 = $396 4 TV's & 1 computer: 4*$72 + 1*$90 = $378 0 TV's & 4 computers: 0*$72 + 4*$90 = $360

OpenStudy (texaschic101):

x = t.v's and y = computers (A) 5(72) + 2(90) = 360 + 180 = 540 3(72) + 3(90) = 216 + 270 = 486 1(72) + 4(90) = 72 + 360 = 432 (B) 4(72) + 0(90) = 288 ---> will not work (C) 3(72) + 1(90) = 216 + 90 = 306 --> will not work (D) 4(72) + 0 (90) = 288 --> will not work This could be completely wrong....but I am going to say A.